# Suppression of Bogoliubov momentum pairing and emergence of non-Gaussian correlations in ultracold interacting Bose gases

### v1

@@ Abstract
When interactions become strong enough as to push the system away from the weakly interacting regime, the reconstruction of correlations going beyond those predicted by Bogoliubov theory becomes a challenge, for both experiments and theory. On the theory side, the main theoretical tool to treat strong correlations is offered by numerical calculations Ceperley (1995); Svistunov et al. (2015) – which nonetheless rarely address the correlations probing non-Gaussian statistics, because of their significant numerical cost. On the experimental front, ultracold atoms offer the possibility to control the strength of interactions, giving access to the rich showcase of quantum many-body phenomena exhibited by strongly interacting bosons Fisher et al. (1989); Bloch et al. (2008); Georgescu et al. (2014). Resonant two-body interactions tuned with Feshbach resonances realize unitary Bose gases Chevy and Salomon (2016) hosting few-body Efimov physics Makotyn et al. (2014); Fletcher et al. (2017); Eismann et al. (2016); Klauss et al. (2017); Eigen et al. (2018); Yan et al. (2020), which are nonetheless plagued by large three-body losses Chevy and Salomon (2016) and observing their equilibrium behavior is challenging. But strongly-interacting Bose gases are also realized in optical lattices, where three-body losses are negligible, offering a clear path to study equilibrium many-body physics. Moreover the use of detection methods with single-atom resolution, as developed in our experiment using metastable helium-4 (${}^{4}$$*{}^{*}$) atoms Cayla et al. (2018), is instrumental in reconstructing correlations to all orders Hercé et al. (2023).
@@
Here the sums are restricted to momenta ${\bm{k}},{\bm{k}^{\prime}},{\bm{k}^{\prime\prime}}$ such that only operators of modes with non-zero momentum are included. Bogoliubov theory, mentioned in the introduction, amounts to retaining only the quadratic term Eq. (1) in the above expression, which corresponds to linearizing quantum fluctuations, and describing only pairwise correlations among opposite momenta.
@@
Figure 2: Momentum density profiles and two-body connected correlations at opposite momenta. a Normalized 1D cuts of the momentum density $\rho(k)/\rho(0)$ for varying ratios $u=U/J\in[2;22]$. b Same as panel a in semi-log scale. The green shaded area represents the integration volume $\Omega_{k}$ of the depletion over which two-body correlations are computed. c Condensate fraction $f_{c}$ as a function of $u=U/J$. The dashed line is a guide to the eye $\propto(1-u/u_{c})^{2\beta}$ with $\beta=0.3485$ and $u_{c}=26$ Hercé et al. (2021). Error bars reflect one standard deviation. d Amplitude $\overline{G}_{\rm c}^{(2)}(\bm{K}=\bm{0})$ of the Bogoliubov momentum pairing at opposite momenta (${\bm{K}}={\bm{k}}_{1}+{\bm{k}}_{2}={\bm{0}}$) measured in the volume $\Omega_{k}$ occupied by the depletion of the lattice BEC (orange circles) and of a harmonically-trapped BEC (orange square). The dashed-line is a guide to the eye. The vertical error bars correspond to $68\%$-confidence intervals obtained from a bootstrap method. Inset: $\overline{G}_{\rm c}^{(2)}(\bm{K}=\bm{0})$ plotted as a fonction of the condensate fraction $f_{c}$.
@@
The emergence of correlated sets of $n>2$ modes, whose correlations are not reducible to pairwise ones, clearly brings to a departure from the Gaussian-state description. A quantitative understanding of this regime, and in particular of the structure of its correlations, remains challenging (see e.g. Ref. Carlen et al. (2021) for a recent theoretical effort in continuum space). To address this problem experimentally, we probe correlated sets of $n$ momentum modes (depicted in Fig. 1a-c) via connected correlation functions $G_{\rm c}^{(n)}$ which are $n$-th order multivariate cumulants Ursell (1927). Here, we probe connected correlations among $n$ momentum populations, $G_{\rm c}^{(n)}(N_{{\bm{k}}_{1}},N_{{\bm{k}}_{2}},...,N_{{\bm{k}}_{n}})=\langle N_{{\bm{k}}_{1}}N_{{\bm{k}}_{2}}...N_{{\bm{k}}_{n}}\rangle-G_{\rm dis}^{(n)}(N_{{\bm{k}}_{1}},N_{{\bm{k}}_{2}},...,N_{{\bm{k}}_{n}})$ where the disconnected part $G_{\rm dis}^{(n)}$ includes the contributions of all lower-order correlations $n^{\prime}<n$ to the total $n$-body correlations. The $G_{\rm c}^{(n)}$ function isolates therefore genuine $n$-mode correlations, i.e. those not reducible to correlations of lower order. Connected correlations among zero-sum momentum modes, $G_{\rm c}^{(n)}(\bm{K}=0)$, represent the most relevant form of quantum correlations induced by interactions, both at the level of Bogoliubov theory as well as beyond (Fig. 1a-c). On the other hand, incoherent (thermal) fluctuations are not subject to any momentum-sum constraint (as the momentum sum itself can fluctuate incoherently). Hence detecting correlations among zero-sum modes is a strong indication of their quantum origin, caused by processes of the kind depicted in Fig. 1a-c. In the following, we shall focus our attention on the two-mode, zero-sum connected correlations $G_{\rm c}^{(2)}(\bm{k},-\bm{k})=\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle-\langle N_{\bm{k}}\rangle\langle N_{-\bm{k}}\rangle$. Bogoliubov theory predicts these correlations to take the form $G^{(2)}_{c}({\bm{k}},-{\bm{k}})=|\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}\rangle|^{2}$ (Supplementary Information), namely as stemming from momentum pairing in the system.
@@
Our experiment begins with the production of a metastable helium-4 $\left({}^{4}\textrm{He}^{*}\right)$ Bose-Einstein condensate (BEC) in a crossed optical dipole trap Bouton et al. (2015). The BEC is then adiabatically transferred to a 3D optical lattice Carcy et al. (2021), which we use to increase the interactions and the quantum depletion Tenart et al. (2021). In the lattice, the Bose gas is well described by the Bose-Hubbard Hamiltonian, $H=-J\sum_{\langle ij\rangle}(b_{i}^{\dagger}b_{j}+b_{j}^{\dagger}b_{i})+V_{\rm int}$ with tunneling energy $J$ and on-site interaction energy $U$ – here $\langle ij\rangle$ denotes a pair of nearest neighbors and $b_{j}^{\dagger}$ (resp. $b_{j}$) the creation (resp. destruction) operator on site $j$. The lattice wave vector is $k_{d}=2\pi/d$ where the $d=775$ nm is the lattice spacing. We access momentum correlations in a time-of-flight (TOF) experiment where the ${}^{4}$$*{}^{*}$ atoms fall onto a pair of Micro-Channel Plates (MCPs) located 434343~{}43cm below the optical lattice. The combination of MCPs with delay lines Nogrette et al. (2015) yields the 3D single-atom-resolved distribution of the gas in momentum space, as illustrated in Fig. 1d. We use a total atom number $N=5.0(7)\times 10^{3}$ to ensure that the lattice filling is unity in the center of the trap and that interactions are negligible during the TOF dynamics. Under these conditions, we are able to measure the in-trap momenta of individual atoms Cayla et al. (2018); Tenart et al. (2020).
@@
We have recorded atom distributions at various interaction strengths $u=U/J$ in the range $u\in[2,22]$ ($\sim 2300$ distributions for each $u$), far from the transition to the (finite-temperature) Mott insulator, occurring at the critical ratio $u_{c}=26(1)$ in our experiment Hercé et al. (2021). This is illustrated in Figure 2a, where normalized 1D cuts of the momentum density are shown to exhibit contrasted peaks associated to the presence of large condensate fractions. The semi-logarithmic scale in Figure 2b highlights the increase in the depletion density by more than one order of magnitude at the edge of the first Brillouin zone $k=0.5~{}k_{d}$. The corresponding condensate fraction $f_{c}$ is plotted as a function of interactions $u=U/J$ in Fig. 2c. Importantly, the reduced temperature $T/J$ is constant (within error bars) in the range studied here Carcy et al. (2021). The observed increase of the depletion in Fig. 2a-c thus results predominantly from the increase of interactions (i.e. not an effect of heating). In the following we concentrate our attention on a region of momentum space, denoted by $\Omega_{k}$ (Fig. 2b), which captures the quantum depletion of the condensate. In this work, we use for $\Omega_{k}$ a cubic corona comprised between two cubes centred on ${\bm{k}}={\bm{0}}$ and of sides $2\times 0.7~{}k_{d}$ and $2\times 0.3~{}k_{d}$ respectively (Fig. 2b and Methods).
@@
From the experimental data, we extract the connected correlations at opposite momenta, $G_{\rm c}^{(2)}(\bm{k},-\bm{k})$, for momenta $\bm{k}$ belonging to $\Omega_{k}$. To increase the signal-to-noise ratio, we sum $G_{\rm c}^{(2)}(\bm{k},-\bm{k})$ over the modes occupying the volume $\Omega_{k}$, defining the integrated amplitude $\overline{G}_{\rm c}^{(2)}(\bm{0})=\sum_{\bm{k}\in\Omega_{k}}\ G_{\rm c}^{(2)}(\bm{k},-\bm{k})$. In Figure 2d, we plot $\overline{G}_{\rm c}^{(2)}(\bm{0})$ as a function of the interaction strength $u$. The measured values have been multiplied by $\eta^{-2}$ to account for the finite detection efficiency $\eta=0.53(2)$ of our detector. At small values of $u$, $\overline{G}_{\rm c}^{(2)}(\bm{0})$ increases with $u$. This observation is compatible with Bogoliubov theory, which predicts that the two-mode correlated quantum depletion always increases with $u$ (Fig. 3a) Toth et al. (2008). In contrast, our measurement shows that $\overline{G}_{\rm c}^{(2)}(\bm{0})$ decreases at larger values of $u$. The position of the maximum corresponds to a relatively large condensate fraction of $f_{c}\sim 0.75$ (inset of Fig. 2d), far from the Mott transition. This strongly suggests that our observations are not specific to lattice bosons at integer filling, but are generic of strongly interacting Bose gases.
@@
We stress that this non-monotonic variation of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ results from an adiabatic change of the interaction strength Carcy et al. (2021), and is not the result of heating. It should also be contrasted with the variation of the bunching amplitude $g^{(2)}_{\rm HBT}(0)=\left(\sum_{\bm{k}\in\Omega_{k}}\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{\bm{k}}a_{\bm{k}}a_{\bm{k}}\rangle\right)/\left(\sum_{\bm{k}\in\Omega_{k}}\langle N_{\bm{k}}\rangle^{2}\right)$ set by the Bose statistics. Indeed, we find a constant and perfectly-contrasted value $g^{(2)}_{\rm HBT}(0)=2.00(5)$ in all regimes of interactions (Supplementary Information). This result, expected at weak interactions, extends our previous observations Cayla et al. (2020) to the strongly-interacting regime. Finally, a further observation is the increase of the fluctuations of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ with interactions shown by the error bars (see Fig. 2d) obtained from a bootstrap analysis of the connected correlations measured in each experimental shot. This analysis suggests that the number of pairs at opposite momenta fluctuates more as interactions increase. An interesting question for a future work is to understand whether this behaviour stems from the increasing complexity of the equilibrium many-body state at strong interactions, made of correlated subsets with $n>2$ modes.
@@
To investigate the fate of two-mode connected correlations $\overline{G}_{\rm c}^{(2)}(\bm{0})$ beyond the weakly interacting regime theoretically, we make use of numerically exact quantum Monte Carlo (QMC) simulations on a 3D quantum rotor model (Supplementary Information), which captures the physics of the Bose-Hubbard model at large, integer filling $n_{\rm QR}$, and without confining potentials Wallin et al. (1994). Calculations on a quantum-rotor model are not expected to be quantitatively accurate when compared with experiments conducted at nearly unit filling in a harmonic trap. Yet, unlike QMC simulations for the Bose-Hubbard model, they give access to arbitrary correlations in momentum space without the need for specialized update algorithms Roscilde (2008); Fang et al. (2016); and they do so across all regimes of temperatures and interactions up to the Mott transition (Supplementary Information). Fig. 3a shows $\overline{G}_{\rm c}^{(2)}(\bm{0})$, integrated over quasi-momenta in a cubic corona $\Omega_{k}$ between two cubes of sides $2\times 0.3k_{d}$ and $2\times 0.5k_{d}$, for a system of $10^{3}$ cubic lattice sites, as a function of the interaction strength, and for different temperatures. The theoretical data have been normalized to the atom number $\langle N_{\Omega_{k}}\rangle$ detected experimentally at a given interaction strength $u$ (Methods). The temperature of the experiment was previously estimated to be around $k_{B}T/J\approx 2$ Carcy et al. (2021), corresponding to the data for the quantum rotor model at a temperature $T_{J}=k_{B}T/(2Jn_{\rm QR})=1$.
@@
Figure 3: Numerical calculations of connected correlations $\overline{G}_{\rm c}^{(n)}(\bm{0})$ with the quantum rotor model. a Two-mode connected correlations $\overline{G}_{\rm c}^{(2)}(\bm{0})$ obtained with the quantum rotor model (solid lines) at different temperatures, ranging from $T_{J}=k_{B}T/2n_{\rm QR}J=0.2$ (dark blue) to $T_{J}=1.4$ (dark red) in steps of 0.4. The system size is $10^{3}$. The dotted line is the result from Bogoliubov theory for the Bose-Hubbard model with $10^{3}$ sites and unity filling. Inset: log-plot of the data from the main panel. The amplitude of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ is normalized to the measured atom number $\langle N_{\Omega_{k}}\rangle_{\rm exp}$ in the experiment (Methods and Supplementary Information). b Three-mode connected correlations $\overline{G}_{\rm c}^{(3)}(\bm{0})$ obtained with the quantum rotor model at the various temperatures. The amplitude of $\overline{G}_{\rm c}^{(3)}(\bm{0})$ is normalized to the measured atom number (Methods and Supplementary Information).
@@
As we shall now explain, measuring $\overline{G}_{\rm c}^{(3)}(0)\neq 0$ is, however, not the only means to demonstrate non-Gaussian statistics. In the spirit of Bogoliubov theory, we relax the assumption of particle-number conservation, but we consider that all the interaction processes described in Eqs. (1), (2) and (3) do not lead to the appearance of a finite value of $\langle a_{\bm{k}}\rangle$ in the depletion. This assumption is compatible with the perfectly-contrasted bunching, $g^{(2)}_{\rm HBT}(0)=2.00(5)$, measured in the experiment as a finite value of $\langle a_{\bm{k}}\rangle$ would reduce the bunching amplitude. Hence, the connected two-mode correlations can then be expressed in terms of the 4-th-order cumulant $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}$ (Supplemental Material),
$G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ $=$ $|\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}\rangle|^{2}+|\langle a^{\dagger}_{\bm{k}}a_{-\bm{k}}\rangle|^{2}$ (4)
Eq. (4) shows that negative values of $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ necessarily imply that $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}\neq 0$, which is a signature of non-Gaussian statistics. Indeed a negative correlation $\overline{G}_{\rm c}^{(2)}({\bm{k}},-{\bm{k}})<0$ amounts to anti-bunching, namely the fact that the coincidence counts $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle$ is smaller than what expected in the absence of correlations, $\langle N_{\bm{k}}\rangle\langle N_{-\bm{k}}\rangle$. This only occurs in the presence of connected correlations of order $n>2$.
In the experiment, we observe a drastic change in the shape of the integrated correlation function $\overline{G}_{\rm c}^{(2)}({\bm{k}},-{\bm{k}}+\delta{\bm{k}})$ at large interactions $u=20$ and $22$: there is hardly any correlation peak at $\delta{\bm{k}}={\bm{0}}$ but a constant (negative) offset as a function of $\delta{\bm{k}}$ (Supplemental Material). In this regime, we thus find $\overline{G}_{\rm c}^{(2)}(\bm{0})<0$ (see Fig. 2d) and $\sum_{{\bm{k}}}\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}\neq 0$. This non-zero fourth-order cumulant is an unambiguous signature of non-Gaussian statistics that does not rely on any theoretical modeling. As such, it represents a striking evidence of the strongly correlated (non-Gaussian) nature of Bose gases at large interactions.

### v2

@@ Abstract
When interactions become strong enough as to push the system away from the weakly interacting regime, the reconstruction of correlations going beyond those predicted by Bogoliubov theory becomes a challenge, for both experiments and theory. On the theory side, the main theoretical tool to treat strong correlations is offered by numerical calculations Ceperley (1995); Svistunov et al. (2015) – which nonetheless rarely address the correlations probing non-Gaussian statistics, because of their significant numerical cost. On the experimental front, ultracold atoms offer the possibility to control the strength of interactions, giving access to the rich showcase of quantum many-body phenomena exhibited by strongly interacting bosons Fisher et al. (1989); Bloch et al. (2008); Georgescu et al. (2014). Resonant two-body interactions tuned with Feshbach resonances realize unitary Bose gases Chevy and Salomon (2016) hosting few-body Efimov physics Makotyn et al. (2014); Fletcher et al. (2017); Eismann et al. (2016); Klauss et al. (2017); Eigen et al. (2018); Yan et al. (2020), which are nonetheless plagued by large three-body losses Chevy and Salomon (2016) and observing their equilibrium behavior is challenging. But strongly-interacting Bose gases are also realized in optical lattices, where three-body losses are negligible, offering a clear path to study equilibrium many-body physics. Moreover the use of detection methods with single-atom resolution, as developed using metastable helium-4 (4He) atoms Schellekens et al. (2005); Vassen et al. (2012); Cayla et al. (2018), is instrumental in reconstructing correlations to all orders Dall et al. (2013); Hercé et al. (2023).
@@
Here the sums are restricted to momenta ${\bm{k}},{\bm{k}^{\prime}},{\bm{k}^{\prime\prime}}$ such that only operators of modes with non-zero momentum are included. Bogoliubov theory, mentioned in the introduction, amounts to retaining only the quadratic term of Eq. (1) in the above expression, which corresponds to linearizing quantum fluctuations, and describing only pairwise correlations among opposite momenta.
@@
The emergence of correlated sets of $n>2$ modes, whose correlations are not reducible to pairwise ones, distinctly indicates a deviation from the Gaussian-state depiction. A quantitative understanding of this regime, and in particular of the structure of its correlations, remains challenging (see e.g. Ref. Carlen et al. (2021) for a recent theoretical effort in continuum space). To address this problem experimentally, we probe correlated sets of $n$ momentum modes (depicted in Fig. 1a-c) via connected correlation functions $G_{\rm c}^{(n)}$ which are $n$-th order multivariate cumulants Ursell (1927). Here, we probe connected correlations among $n$ momentum populations, $G_{\rm c}^{(n)}(N_{{\bm{k}}_{1}},N_{{\bm{k}}_{2}},...,N_{{\bm{k}}_{n}})=\langle N_{{\bm{k}}_{1}}N_{{\bm{k}}_{2}}...N_{{\bm{k}}_{n}}\rangle-G_{\rm dis}^{(n)}(N_{{\bm{k}}_{1}},N_{{\bm{k}}_{2}},...,N_{{\bm{k}}_{n}})$ where the disconnected part $G_{\rm dis}^{(n)}$ includes the contributions of all lower-order correlations $n^{\prime}<n$ to the total $n$-body correlations. The $G_{\rm c}^{(n)}$ function isolates therefore genuine $n$-mode correlations, i.e. those not reducible to correlations of lower order. Connected correlations among zero-sum momentum modes, $G_{\rm c}^{(n)}(\bm{K}=0)$, represent the most relevant manifestation of quantum fluctuations induced by interactions, both at the level of Bogoliubov theory as well as beyond (Fig. 1a-c). In contrast, incoherent (thermal) fluctuations are not subject to any momentum-sum constraint (as the momentum sum itself can fluctuate incoherently). Hence detecting correlations among zero-sum modes is a strong indication of their quantum origin, caused by processes of the kind depicted in Fig. 1a-c. In the following, we shall focus our attention on the two-mode, zero-sum connected correlations $G_{\rm c}^{(2)}(\bm{k},-\bm{k})=\langle N_{\bm{k}}N_{-\bm{k}}\rangle-\langle N_{\bm{k}}\rangle\langle N_{-\bm{k}}\rangle$. Bogoliubov theory predicts these correlations to take the form $G^{(2)}_{c}({\bm{k}},-{\bm{k}})=|\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}\rangle|^{2}$ (Supplementary Information), namely as stemming from momentum pairing in the system.
@@
Our experiment begins with the production of a metastable helium-4 $\left({}^{4}\textrm{He}^{*}\right)$ Bose-Einstein condensate (BEC) in a crossed optical dipole trap Bouton et al. (2015). The BEC is then adiabatically transferred to a 3D optical lattice Carcy et al. (2021), which we use to increase the interactions and the quantum depletion Tenart et al. (2021). In the lattice, the Bose gas is well described by the Bose-Hubbard Hamiltonian, $H=-J\sum_{\langle ij\rangle}(b_{i}^{\dagger}b_{j}+b_{j}^{\dagger}b_{i})+V_{\rm int}$ with tunneling energy $J$ and on-site interaction energy $U$ – here $\langle ij\rangle$ denotes a pair of nearest neighbors and $b_{j}^{\dagger}$ (resp. $b_{j}$) the creation (resp. destruction) operator on site $j$. The lattice wave vector is $k_{d}=2\pi/d$ where the $d=775$ nm is the lattice spacing. We access momentum correlations in a time-of-flight (TOF) experiment where the 4He atoms fall onto a pair of Micro-Channel Plates (MCPs) located 434343~{}43cm below the optical lattice. The combination of MCPs with delay lines Nogrette et al. (2015) yields the 3D single-atom-resolved distribution of the gas in momentum space, as illustrated in Fig. 1d. We use a total atom number $N=5.0(7)\times 10^{3}$ to ensure that the lattice filling is unity in the center of the trap and that interactions are negligible during the TOF dynamics. Under these conditions, we are able to measure the in-trap momenta of individual atoms Cayla et al. (2018); Tenart et al. (2020).
Figure 2: Momentum density profiles and two-body connected correlations at opposite momenta. a Normalised 1D cuts of the momentum density $\rho(k)/\rho(0)$ for varying ratios $u=U/J\in[2;22]$. b Same as panel a in semi-log scale. The green shaded area represents the integration volume $\Omega_{k}$ of the depletion over which two-body correlations are computed. c Condensate fraction $f_{c}$ as a function of $u=U/J$. The dashed line is a guide to the eye $\propto(1-u/u_{c})^{2\beta}$ with $\beta=0.3485$ and $u_{c}=26$ Hercé et al. (2021). Error bars reflect one standard deviation. d Amplitude $\overline{G}_{\rm c}^{(2)}(\bm{K}=\bm{0})$ of the Bogoliubov momentum pairing at opposite momenta (${\bm{K}}={\bm{k}}_{1}+{\bm{k}}_{2}={\bm{0}}$) measured in the volume $\Omega_{k}$ occupied by the depletion of the lattice BEC (orange circles). The measured values have been multiplied by $\eta^{-2}$ to account for the finite detection efficiency $\eta=0.53(3)$ of our detector. The vertical error bars correspond to $68\%$-confidence intervals obtained from a bootstrap method. Inset: $\overline{G}_{\rm c}^{(2)}(\bm{K}=\bm{0})$ plotted as a function of the condensate fraction $f_{c}$.
@@
We have recorded atom distributions at various interaction strengths $u=U/J$ in the range $u\in[1,22]$ ($\sim 2300$ distributions for each $u$), far from the transition to the (finite-temperature) Mott insulator, occurring at the critical ratio $u_{c}=26(1)$ in our experiment Hercé et al. (2021). This is illustrated in Figure 2a, where normalised 1D cuts of the momentum density are shown to exhibit contrasted peaks associated to the presence of large condensate fractions. The semi-logarithmic scale in Figure 2b highlights the increase in the depletion density by more than one order of magnitude at the edge of the first Brillouin zone $k=0.5~{}k_{d}$. The corresponding condensate fraction $f_{c}$ is plotted as a function of interactions $u=U/J$ in Fig. 2c. Importantly, we perform a thermometry Carcy et al. (2021) and find that the change in the reduced temperature $T/J$ is compatible with an adiabatic variation with $U/J$ (Supplementary Information). The observed increase of the depletion in Fig. 2a-c thus results predominantly from the increase of interactions (i.e. not an effect of heating). In the following we concentrate our attention on a region of momentum space, denoted by $\Omega_{k}$ (Fig. 2b), which captures the quantum depletion of the condensate. In this work, we use for $\Omega_{k}$ a cubic corona comprised between two cubes centred on ${\bm{k}}={\bm{0}}$ and of sides $2\times 0.7~{}k_{d}$ and $2\times 0.3~{}k_{d}$ respectively (Fig. 2b and Methods).
@@
From the experimental data, we extract the connected correlations at opposite momenta, $G_{\rm c}^{(2)}(\bm{k},-\bm{k})$, for momenta $\bm{k}$ belonging to $\Omega_{k}$. To increase the signal-to-noise ratio, we sum $G_{\rm c}^{(2)}(\bm{k},-\bm{k})$ over $\Omega_{k}$, defining the integrated amplitude $\overline{G}_{\rm c}^{(2)}(\bm{0})=\sum_{\bm{k}\in\Omega_{k}}\ G_{\rm c}^{(2)}(\bm{k},-\bm{k})$. In Figure 2d, we plot $\overline{G}_{\rm c}^{(2)}(\bm{0})$ as a function of the interaction strength $u$. We find that $\overline{G}_{\rm c}^{(2)}(\bm{0})\approx 0$ as $u\rightarrow 0$, confirming the absence of momentum pairing in non-interacting gases. Furthermore, $\overline{G}_{\rm c}^{(2)}(\bm{0})$ first rises as $u$ is increased from $u=0$. This observation is an excellent agreement with Bogoliubov theory, which predicts that the two-mode correlated quantum depletion always increases with $u$ (Fig. 3a) Toth et al. (2008). In contrast, our measurement shows that $\overline{G}_{\rm c}^{(2)}(\bm{0})$ decreases at larger values of $u$. The position of the maximum corresponds to a relatively large condensate fraction of $f_{c}\sim 0.75$ (inset of Fig. 2d), far from the Mott transition. This strongly suggests that our observations are not specific to lattice bosons at integer filling, but are generic of strongly interacting Bose gases.
@@
We stress that this non-monotonic variation of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ results from an adiabatic change of the interaction strength Carcy et al. (2021), and is not the result of heating. It should also be contrasted with the variation of the bunching amplitude $g^{(2)}_{\rm HBT}(0)=\left(\sum_{\bm{k}\in\Omega_{k}}\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{\bm{k}}a_{\bm{k}}a_{\bm{k}}\rangle\right)/\left(\sum_{\bm{k}\in\Omega_{k}}\langle N_{\bm{k}}\rangle^{2}\right)$ set by the Bose statistics. Indeed, we find a constant and perfectly-contrasted value $g^{(2)}_{\rm HBT}(0)=2.00(5)$ in all regimes of interactions (Supplementary Information). This result, expected at weak interactions, extends our previous observations Cayla et al. (2020) to the strongly-interacting regime. Finally, a further observation is the increase of the fluctuations of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ with interactions shown by the error bars (see Fig. 2d) obtained from a bootstrap analysis of the measured connected correlations. This analysis suggests that the number of pairs at opposite momenta fluctuates more as interactions increase. An interesting question for a future work is to understand whether this behaviour stems from the increasing complexity of the equilibrium many-body state at strong interactions, made of correlated subsets with $n>2$ modes.
@@
To investigate the fate of two-mode connected correlations $\overline{G}_{\rm c}^{(2)}(\bm{0})$ beyond the weakly interacting regime theoretically, we make use of numerically exact quantum Monte Carlo (QMC) simulations on a 3D quantum rotor model (Supplementary Information), which captures the physics of the Bose-Hubbard model at large, integer filling $n_{\rm QR}$, and without confining potentials Wallin et al. (1994). Calculations on a quantum-rotor model are not expected to be quantitatively accurate when compared with experiments conducted at nearly unit filling in a harmonic trap. Yet, unlike QMC simulations for the Bose-Hubbard model, they give access to arbitrary correlations in momentum space without the need for specialized update algorithms Roscilde (2008); Fang et al. (2016); and they do so across all regimes of temperatures and interactions up to the Mott transition (Supplementary Information). Fig. 3a shows $\overline{G}_{\rm c}^{(2)}(\bm{0})$, integrated over quasi-momenta in a cubic corona $\Omega_{k}$ between two cubes of sides $2\times 0.3k_{d}$ and $2\times 0.5k_{d}$, for a system of $10^{3}$ cubic lattice sites, as a function of the interaction strength, and for different temperatures. The theoretical data have been normalised to the atom number $\langle N_{\Omega_{k}}\rangle$ detected experimentally at a given interaction strength $u$ (Methods). The temperature of the experiment was previously estimated to be around $k_{B}T/J\approx 2$ Carcy et al. (2021), corresponding to the data for the quantum rotor model at a temperature $T_{J}=k_{B}T/(2Jn_{\rm QR})=1$.
@@
Figure 3: Numerical calculations of connected correlations $\overline{G}_{\rm c}^{(n)}(\bm{0})$ with the quantum rotor model. a Two-mode connected correlations $\overline{G}_{\rm c}^{(2)}(\bm{0})$ obtained with the quantum rotor model (solid lines) at different temperatures, ranging from $T_{J}=k_{B}T/2n_{\rm QR}J=0.2$ (dark blue) to $T_{J}=1.4$ (dark red) in steps of 0.4. The system size is $10^{3}$. The dotted line is the result from Bogoliubov theory for the Bose-Hubbard model with $10^{3}$ sites and unity filling. Inset: log-plot of the data from the main panel. The amplitude of $\overline{G}_{\rm c}^{(2)}(\bm{0})$ is normalised to the measured atom number $\langle N_{\Omega_{k}}\rangle_{\rm exp}$ in the experiment (Methods and Supplementary Information). b Three-mode connected correlations $\overline{G}_{\rm c}^{(3)}(\bm{0})$ obtained with the quantum rotor model at the various temperatures. The amplitude of $\overline{G}_{\rm c}^{(3)}(\bm{0})$ is normalised to the measured atom number (Methods and Supplementary Information). The shaded area represents the uncertainty of one standard deviation.
@@
As we shall now explain, measuring $\overline{G}_{\rm c}^{(3)}(0)\neq 0$ is, however, not the only means to demonstrate non-Gaussian statistics. In the spirit of Bogoliubov theory, we relax the assumption of particle-number conservation, but we consider that all the interaction processes described in Eqs. (1)-(3) do not lead to the appearance of a finite value of $\langle a_{\bm{k}}\rangle$ in the depletion. This assumption is compatible with the perfectly-contrasted bunching, $g^{(2)}_{\rm HBT}(0)=2.00(5)$, measured in the experiment as a finite value of $\langle a_{\bm{k}}\rangle$ would reduce the bunching amplitude. Hence, the connected two-mode correlations can then be expressed in terms of the 4-th-order cumulant $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}$ (Supplemental Material),
$G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ $=$
$|\langle a^{\dagger}_{\bm{k}}$ $a^{\dagger}_{-\bm{k}}\rangle|^{2}+|\langle a^{\dagger}_{\bm{k}}a_{-\bm{k}}\rangle|^{2}+\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}.$ (4)
Equation (4) shows that negative values of $G^{(2)}_{c}({\bm{k}},-{\bm{k}})$ necessarily imply that $\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}\neq 0$, which is a signature of non-Gaussian statistics. A similar approach to unveil a non-zero 4-th-order cumulant from the measurement of two-photon correlations has been used recently Ferioli et al. (2024). A negative correlation $G_{\rm c}^{(2)}({\bm{k}},-{\bm{k}})<0$ amounts to anti-bunching, namely the fact that the coincidence counts $\langle N_{\bm{k}}N_{-\bm{k}}\rangle$ is smaller than what expected in the absence of correlations, $\langle N_{\bm{k}}\rangle\langle N_{-\bm{k}}\rangle$. Our numerical results on quantum rotors exhibit $G_{\rm c}^{(2)}({\bm{k}},-{\bm{k}})<0$ at large values $U/J$, when the homogeneous system enters the normal phase (Supplemental Material). This stems from the conservation of the total atom number, which indeed corresponds to non-Gaussian correlations as it involves all atoms.
In the experiment, we find a drastic change in the shape of $\overline{G}_{\rm c}^{(2)}(\delta{\bm{k}})$ at large interactions $u\geq 20$: there is hardly any correlation peak at $\delta{\bm{k}}={\bm{0}}$ but a constant (negative) offset independent of $\delta{\bm{k}}$ (Supplemental Material). This anti-bunching thus differs from that of two fermions which takes place at short separations Rom et al. (2006): the absence of a characteristic two-body length scale suggests that the observed $\overline{G}_{\rm c}^{(2)}({\bm{0}})<0$ is set by correlations between more than two atoms in momentum space. As discussed above, $\overline{G}_{\rm c}^{(2)}(\bm{0})<0$ implies $\sum_{{\bm{k}}}\langle a^{\dagger}_{\bm{k}}a^{\dagger}_{-\bm{k}}a_{\bm{k}}a_{-\bm{k}}\rangle_{c}\neq 0$, which is an unambiguous signature of non-Gaussian statistics. Importantly, this signature does not rely on any theoretical modeling. As such, it represents a striking evidence of the strongly correlated (non-Gaussian) nature of Bose gases at large interactions. Determining whether the observed $\overline{G}_{\rm c}^{(2)}({\bm{0}})<0$ in the harmonically-trapped gas shares the same physical origin as that found in numerical studies of homogeneous systems is a complex question that requires further investigation.